OpenStudy (anonymous):

Find three solutions of the equation. y= 3x+4 a. (-2,-2), (1,7), (2,11) b. (0,4), (3,11), (-2,-2) c. (-2,-2), (1,7), (0,4) d. (1,7), (0,4), (-1,0)

OpenStudy (anonymous):

@tyler**collins

OpenStudy (jordanloveangel):

one sec

OpenStudy (anonymous):

okay (:

OpenStudy (jordanloveangel):

Simplifying y = 3x + 4 Reorder the terms: y = 4 + 3x Solving y = 4 + 3x Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Simplifying y = 4 + 3x

OpenStudy (anonymous):

You could really just plug all in to get them, but Ill make a graph if Jordan doesn't

OpenStudy (jordanloveangel):

okay thats the best i can do

OpenStudy (anonymous):

We know y=mx+b form will give you all the solutions on a certain equation.

OpenStudy (anonymous):

As long as all three coordinates lay on that line it should be correct

OpenStudy (anonymous):

(0,4) is one

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

Is (1,7) and (-1,0)

OpenStudy (anonymous):

yess thank you (:

OpenStudy (anonymous):

Welcomes ^-^